need son hlp on this one
Given f(x)=8(2x^3-9)^7 find the inverse function
thank you

Answer by Nate(3500)

y = 8(2x^3 - 9)^7
x = 8(2y^3 - 9)^7
x/8 = (2y^3 - 9)^7
(x/8)^(1/7) = 2y^3 - 9
(x/8)^(1/7) + 9 = 2y^3
(x/8)^(1/7)/2 + 9/2 = y^3
((x/8)^(1/7)/2 + 9/2)^(1/3) = y
f(x) = 8(2x^3 - 9)^7
f^-1(x) = ((x/8)^(1/7)/2 + 9/2)^(1/3)